Show that e^x Cosy is Harmonic, find its Conjugate harmonic function.

1
Expert's answer
2022-01-12T08:35:45-0500
"u=e^{x}\cos y"

"\dfrac{\partial u}{\partial x}=e^{x}\cos y"

"\dfrac{\partial^2 u}{\partial x^2}=e^{x}\cos y"

"\dfrac{\partial u}{\partial y}=-e^{x}\sin y"

"\dfrac{\partial^2 u}{\partial y^2}=-e^{x}\cos y"


"\dfrac{\partial^2 u}{\partial x^2}+\dfrac{\partial^2 u}{\partial y^2}=e^{x}\cos y-e^{x}\cos y"

Hence


"\nabla^2u=\dfrac{\partial^2 u}{\partial x^2}+\dfrac{\partial^2 u}{\partial y^2}=0, x, y\in \R"

Therefore the function "u=e^{x}\cos y" is harmonic.


We find the harmonic conjugate of "u," denoted by "v," which satisfies the Cauchy-Riemann equations. It follows from 


"\dfrac{\partial u}{\partial x}=e^{x}\cos y=\dfrac{\partial v}{\partial y}"

that "v(x, y) = −e ^x \cos y + φ(x)," where "\u03c6(x)" is a differentiable real-valued function of "x." Also, from


"\dfrac{\partial u}{\partial y}=-e^{x}\sin y=\dfrac{\partial v}{\partial x}"

we have "e ^x \cos y = e ^x \cos y −φ'(x)=>φ(x)" i.e., "\u03c6(x)" is constant. 

Therefore "v(x, y) = −e ^x \cos y + c, c\in \Complex," is the harmonic conjugate of "u."


Hence


"f(z)=u+iv=e ^x(\cos y+i \sin y)+ic"

"=e^{x+iy} +ic=e^z+ic"